The speed of cars on the Laurentian Autoroute are known to form a normal distribution with u=105 km/hr and o= 10 km/hr. If a car was randomly selected, what is the probability that its speed was between 95 and 110 km/hr? Express your answer as a percentage (rounder to two decimal places)
Answer #1

Solution :

Given that ,

mean = = 105

standard deviation = = 10

P(95 < x < 110) = P((95 - 105)/ 10) < (x - ) / < (110 - 105) / 10) )

= P(-1 < z < 0.5)

= P(z < 0.5) - P(z < -1)

= 0.6915 - 0.1587

= 0.5328

= 53.28%

Probability = 53.28 %